# LS-EEND: Long-Form Streaming End-to-End Neural Diarization with Online Attractor Extraction

### v1

@@ I Introduction
Offline diarization systems are effective for applications like meeting minutes. However, in some online scenarios, such as multi-party human-robot interaction or real-time subtitling of video conferences, the system needs to process audio streams and recognize/respond to the speaking person in real time. To meet these demands, several online/streaming diarization methods have been proposed [18, 19, 20, 21, 22, 23, 24, 25, 26]. In an online cascaded system, each module needs to operate in real time. Typically, sufficient audio frames are collected within a chunk (e.g., 2 s) for embedding extraction, and a short chunk shift (e.g., 0.5 s) is used to achieve low diarization latency [18, 19, 20, 21], then the system performs incremental clustering of extracted embeddings. Along the time increases, re-clustering from scratch for each new segment becomes time-consuming. One solution is to cluster with centroids, which tracks global speaker label information [21, 18]. Another approach is to select representative segments for local clustering, which requires addressing the permutation ambiguity problem [19, 20]. For end-to-end models, the offline EEND is adapted for online inference in [25, 26]. A speaker tracing buffer (STB) is introduced to restore previous frames and diarization results. For each new chunk, the new frames are stacked with the buffered frames to perform local diarization. The new diarization results are then permuted according to the buffered diarization results to ensure consistent speaker order. Since all the buffered frames need to be re-fed to the diarization model for each chunk inference, STB-based methods require redundant computations. Additionally, a frame selection strategy needs to be designed carefully to ensure that the buffer is informative enough to solve the speaker permutation ambiguity.
@@ II Related Works
EEND methods integrates speaker embedding extraction and clustering into a single network [10, 11]. To handle flexible number of speakers, Horiguchi et al. [12, 13] propose the EEND-EDA system, where a flexible numbre of attractors are generated by an encoder-decoder-based attractor extractor. Different from the pre-trained unchanging speaker embeddings [30, 31], EEND methods generate changing/local speaker embeddings, namely one speaker is represented with different embeddings when he/she presents in different contexts. Although the extractor can output attractor for unlimited number of speakers, the performance will largely drop when test speakers are more than training speakers. To really handle unlimited number of speakers, [34, 26] proposes to further transform the local attractors obtained within signal blocks to global attractors, then the global attractors can be clustered with an unsupervised clustering method. Note that the number of speakers within one short block is limited, while the number of global speakers can be unlimited.
@@ III-A Label Permutation based on Speaker Appearance Order
In offline scenarios, frame-level speaker embeddings are clustered to determine speaker activity timestamps, where the order of speakers is arbitrary. The diarization model can be trained using the PIT loss [15], similar to techniques employed in the source separation field [35]. However, in streaming scenarios, speakers appear sequentially. The speakers order should thus align with their appearance order.
@@ III-B Model Architecture
Figure 2: Network architecture of the proposed LS-EEND model. ‘$S$+2’ is the number of speakers corresponding to the augmented speaker label in Fig. 1.
@@ III-B1 Causal Embedding Encoder
The causal embedding encoder takes as input the feature sequence $\boldsymbol{X}$ and extracts frame-wise speaker embeddings $\boldsymbol{E}=(\boldsymbol{e}_{1},\ldots,\boldsymbol{e}_{t},\ldots\boldsymbol{e}_{T})$, with $\boldsymbol{e}_{t}\in\mathbb{R}^{D}$ where ${D}$ denotes embedding dimension. As shown in Fig. 2(b), the proposed encoder block employs the architecture of Conformer [36], which has been validated effectiveness for multiple speech processing tasks [37, 38]. The multi-head Retention module is followed by a convolution module then a feed-forward module. Retention and convolution are expected to capture global and local dependencies, respectively. It should be noted that the convolution module in our implementation of Conformer is modified to be causal. $N$ encoder blocks are cascaded.
In most of speaker embedding networks, including our previous work [28], self-attention-based encoders are normally adopted for learning long-term dependency. However, as for streaming diarization, the proposed system needs to continuously process significantly long audio streams (e.g. a meeting of hours), which poses a huge computational challenge due to the quadratic temporal complexity of self-attention. To address this problem, we adapt the Retention mechanism [27] in the embedding encoder to substitute the masked self-attention mechanism [28]. As a successor to the self-attention mechanism, Retention removes the softmax function on the attention values, so that the computation of attention can be formulated in a recursive way and has a linear complexity w.r.t time steps. Similar to self-attention, for a given input sequence $X\in\mathbb{R}^{T\times D}$, Retention also computes the key, query and value vectors, denoted as $K,Q,V$, with trainable matrices $W_{K},W_{Q},W_{V}$, then aggregates the value vectors. However, the specific computations are somehow different. Specifically, Retention can be formulated in a parallel way as:
@@
where and $\odot$ denote matrix/vector transpose and element-wise product, respectively. $\Theta=e^{it\theta}$, $\bar{\Theta}=e^{-it\theta}$ and the exponential decaying term $\Gamma_{t_{1}t_{2}}=\gamma^{t_{1}-t_{2}}$ (for $t_{2}\leq t_{1}$; 0 otherwise) together provides the so-called xPos [39] as a relative position embedding, where $\theta\in\mathbb{R}^{D}$ and $\gamma\in\mathbb{R}$ are constant real numbers. It can be seen that, different from the self-attention mechanism, the softmax function is not applied to the values of $QK^{\top}\odot\Gamma$, instead, a new normalization strategy is proposed to stabilize the numerical scale in [27]. By doing this, Retention can be re-formulated in a recurrent way as:
$\text{Retention}(X_{t})=Q_{t}S_{t},$
@@
where the state matrix $S_{t}$ is recursively updated along time. The parallel representation Eq. (1) is suitable for efficient training, while the recurrent representation Eq. (2) has a linear temporal complexity and can be used for inference. In between, [27] also provides a chuckwise recurrent representation, performing parallel computation within chunk and sate recurrence across chunks, which facilitates the training with long sequences. Retention also uses multiple heads, and multi-scale retention assigns different $\gamma$ for each head as
@@
where $h$ denotes the number of Retention heads. $\gamma$ controls the forgetting rate of past information.
Speaker diarization learns voiceprint of speakers, which are time-invariant in a very long time period, e.g. in the entire audio stream to be processed. In the proposed model, to keep the consistency of speaker embedding along the entire audio stream, we set $\gamma=1$ to not have information decay. In addition, the relative positional embedding terms, i.e. $\Theta$ and $\bar{\Theta}$ in Eq. (1), are removed as positional embedding has no significant impact on the model’s performance, similar to the findings in the previous neural diarization methods [13, 17].
@@ III-B3 Online Attractor Decoder
Decoder Input. At each frame, the speaker embedding is fed into the decoder to generate/update the corresponding attractors. All the $S+2$ attractors share the same repeated embedding as their input source, and the decoder determines whether the embedding is used for attractor updating. To distinguish the input to multiple attractors from each other, we first repeat the speaker embedding $S+2$ times and repeat a positional embedding (PE) [41] along the speaker/attractor index for $T$ times, and then concatenate speaker embedding with PE in the feature dimension. A linear layer is then applied to restore the feature dimension $D$.
Decoder Architecture. To derive the attractor $\boldsymbol{a}_{s,t}$ of the $s$-th speaker at frame $t$, besides the speaker embedding extracted at frame $t$, two more information sources need to be considered: i) the attractors of previous frames for the same speaker, i.e., $\boldsymbol{a}_{s,t^{\prime}},t^{\prime}<t$, which helps maintain the attractor’s consistency of a specific speaker over time. To refer to previous frames, an along-frame Retention module is designed. For the same reason as stated above, in Retention, $\gamma$ is set to 1 and the positional embedding terms are removed. ii) the attractor of other speakers at frame $t$, i.e., $\boldsymbol{a}_{s^{\prime},t},s^{\prime}\neq s$. By leveraging information from other attractors, the distance between attractors can be enlarged to better distinguish different speakers. To this end, we develop a cross-attractor self-attention module, following the standard self-attention mechanism [41]. After, a feed-forward module is employed. Each of the three modules incorporates a residual connection and layer normalization. $M$ decoder blocks are cascaded.
Decoder Output. The decoder output vectors are then L2-norm normalized to have unit length, serving as the final attractors $\boldsymbol{A}_{t}=(\boldsymbol{a}_{0,t},\boldsymbol{a}_{1,t},\ldots,\boldsymbol{a}_{S,t},\boldsymbol{a}_{\text{ter},t})$. Speaker activities are calculated by taking the inner product between the frame-wise attractors and the speaker embedding $\boldsymbol{e}_{t}$ as
$\hat{\boldsymbol{y}}_{t}=\sigma(\boldsymbol{A}_{t}^{\top}\boldsymbol{e}_{t})\in(0,1)^{S+2}$ (4)
@@ III-C Training Loss
where $\mathbb{S}=\{0,1,\ldots,S,{\text{ter}}\}$.
Embedding similarity loss. The clustering effect of the frame-wise speaker embeddings has a salient impact on diarization performance. To regulate the distribution of these embeddings, we introduce an embedding similarity loss. The cosine similarity between each pair of embeddings ($\langle\boldsymbol{e}_{j},\boldsymbol{e}_{k}\rangle,j,k\in\{1,\ldots,T\}$) is constrained to be consistent with the cosine similarity between their corresponding diarization labels ($\langle\boldsymbol{y}_{j},\boldsymbol{y}_{k}\rangle,j,k\in\{1,\ldots,T\}$). The mean square error between the two cosine similarities is minimized as
$\mathcal{L}_{e}=\frac{1}{T(T-1)/2}\sum_{j=1}^{T}\sum_{k=j+1}^{T}\text{MSE}(\langle\boldsymbol{e}_{j},\boldsymbol{e}_{k}\rangle,\langle\boldsymbol{y}_{j},\boldsymbol{y}_{k}\rangle),$ (6)
[…]
@@ III-E Inference
At inference, the proposed model processes audio streams frame by frame, with the recurrent paradigm of Retention and a linear computational complexity w.r.t frames. The maximum number of speakers $S$ is set to the maximum number appeared in the training set. When the actual number of speakers is smaller than the preseted $S$, it can be detected by the speaker termination marker.
@@ IV-A1 Simulated Mixtures
We follow the recipe presented in [26] to simulate speech mixtures using Switchboard-2 (Phase I & II & III), Switchboard Cellular (Part 1 & 2), and 2004-2008 NIST Speaker Recognition Evaluation (SRE) corpora. Background Noise from MUSAN corpus [42] and Simulated Room Impulse Response [43] are applied to the synthesized mixtures. The sampling rate is 8 kHz. The detailed statistical information of the simulated dataset are provided in Table I. For each of the 1$\sim$8 speakers cases, 100,000 mixtures are generated for training. The overlap ratios between speakers are set quite high, i.e. about 29% to 35%. Audio duration is accordingly increased when more speakers are involved. Besides, for each of the 1$\sim$8 speakers cases, another 500 mixtures are generated for test.

### v2

@@ I Introduction
Offline diarization systems are effective for applications like meeting minutes. However, in some online scenarios, such as multi-party human-robot interaction or real-time subtitling of video conferences, the system needs to process audio streams and recognize/respond to the speaking person in real time. To meet these demands, several online/streaming diarization methods have been proposed [18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. In an online cascaded system, each module needs to operate in real time. Typically, sufficient audio frames are collected within a chunk (e.g., 2 s) for embedding extraction, and a short chunk shift (e.g., 0.5 s) is used to achieve low diarization latency [18, 19, 20, 21], then the system performs incremental clustering of extracted embeddings. Along the time increases, re-clustering from scratch for each new segment becomes time-consuming. One solution is to cluster with centroids, which tracks global speaker label information [21, 18]. Another approach is to select representative segments for local clustering, which requires addressing the permutation ambiguity problem [19, 20]. For end-to-end models, the offline EEND is adapted for online inference in [25, 26, 27]. A speaker tracing buffer (STB) is introduced to restore previous frames and diarization results. For each new chunk, the new frames are stacked with the buffered frames to perform local diarization. The new diarization results are then permuted according to the buffered diarization results to ensure consistent speaker order. Since all the buffered frames need to be re-fed to the diarization model for each chunk inference, STB-based methods require redundant computations. Additionally, a frame selection strategy needs to be designed carefully to ensure that the buffer is informative enough to solve the speaker permutation ambiguity. Recently, online diarization methods based on target-speaker tracking are proposed to perform chunk-wise inference by extracting target-speaker embeddings in real time [28, 29].
@@ II Related Works
EEND methods integrates speaker embedding extraction and clustering into a single network [10, 11]. To handle flexible number of speakers, Horiguchi et al. [12, 13] propose the EEND-EDA system, where a flexible number of attractors are generated by an encoder-decoder-based attractor extractor. Different from the pre-trained unchanging speaker embeddings [33, 34], EEND methods generate changing/local speaker embeddings, namely one speaker is represented with different embeddings when he/she presents in different contexts. Although the extractor can output attractor for unlimited number of speakers, the performance will largely drop when test speakers are more than training speakers. To really handle unlimited number of speakers, [37, 27] proposes to further transform the local attractors obtained within signal blocks to global attractors, then the global attractors can be clustered with an unsupervised clustering method. Note that the number of speakers within one short block is limited, while the number of global speakers can be unlimited.
@@ III-A Label Permutation based on Speaker Appearance Order
In offline scenarios, frame-level speaker embeddings are clustered to determine speaker activity timestamps, where the order of speakers is arbitrary. The diarization model can be trained using the PIT loss [15], similar to techniques employed in the source separation field [38]. However, in streaming scenarios, speakers appear sequentially. The speakers’ order should thus align with their appearance order.
@@ III-B Model Architecture
Figure 2: Network architecture of the proposed LS-EEND model. ‘SS+2’ is the number of speakers corresponding to the augmented speaker label in Fig. 1. The data dimensions are presented in the form of “batch dimension $\times$ dimension of one sample in batch”.
@@ III-B1 Causal Embedding Encoder
The causal embedding encoder takes as input the feature sequence $\boldsymbol{X}$ and extracts frame-wise speaker embeddings $\boldsymbol{E}=(\boldsymbol{e}_{1},\ldots,\boldsymbol{e}_{t},\ldots\boldsymbol{e}_{T})$, with $\boldsymbol{e}_{t}\in\mathbb{R}^{D}$ where D{D} denotes embedding dimension. As shown in Fig. 2(b), the proposed encoder block employs the architecture of Conformer [40], which has been validated effectiveness for multiple speech processing tasks [41, 42]. The multi-head Retention module is followed by a convolution module then a feed-forward module. Retention and convolution are expected to capture global and local dependencies, respectively. It should be noted that the convolution module in our implementation of Conformer is modified to be causal. $N$ encoder blocks are cascaded.
In most of speaker embedding networks, self-attention-based encoders are normally adopted for learning long-term dependency. As for streaming diarization, the system needs to continuously process significantly long audio streams (e.g. a meeting of hours) in an online manner. The masked self-attention mechanism can realize online processing by preventing the use of future information, such as in our previous work [31], but it faces a huge computational challenge due to the quadratic temporal complexity of self-attention. Linear temporal complexity can be achieved by truncating the masked self-attention with a constant truncation length, which however constrains the retrieval of full historical information. In this work, we adapt the Retention mechanism [30], which has a linear temporal complexity and retains the capability of full history retrieval. As a successor to the (masked and causal) self-attention mechanism, Retention removes the softmax function on the attention values, so that the computation of attention can be formulated in a recursive way and has a linear complexity w.r.t time steps. Similar to self-attention, for a given input sequence $X\in\mathbb{R}^{T\times D}$, Retention also computes the key, query and value vectors, denoted as $K,Q,V\in\mathbb{R}^{T\times D}$, with trainable matrices $W_{K},W_{Q},W_{V}\in\mathbb{R}^{D\times D}$, then aggregates the value vectors. However, the specific computations are somehow different. Specifically, Retention can be formulated in a parallel way as:
@@
where and $\odot$ denote matrix/vector transpose and element-wise product, respectively. $\Theta,\bar{\Theta}\in\mathbb{C}^{T\times D/2}$ ($\Theta_{t}=e^{it\theta}$, $\bar{\Theta}_{t}=e^{-it\theta}$) and $\Gamma\in\mathbb{R}^{T\times T}$ together provide the so-called xPos [43] as a relative rotary position embedding, where $i=\sqrt{-1}$ denotes the imaginary unit, $\theta\in\mathbb{R}^{1\times D/2}$ is a constant vector. The relative positional embedding terms, i.e. $\Theta$ and $\bar{\Theta}$ are removed in this work as positional embedding has no significant impact on the model’s performance, similar to the findings in the previous neural diarization methods [13, 17]. $\Gamma$ ($\Gamma_{t_{1}t_{2}}=\gamma^{t_{1}-t_{2}}$, for $t_{2}\leq t_{1}$; 0 otherwise) is a lower triangular mask matrix, which combines causal masking (similar to masked self-attention [44]) and exponential decay, where $\gamma\in\mathbb{R}$ is a constant exponential decay factor. It can be seen that, different from the self-attention mechanism, the softmax function is not applied to the attention scores of $QK^{\top}\odot\Gamma$. By doing this, Retention can be re-formulated in a recurrent way as:
$\text{Retention}(X_{t})=Q_{t}S_{t}$
@@
where $K_{t},Q_{t},V_{t}\in\mathbb{R}^{1\times D}$ are the key, query and value vectors at frame $t$, $S_{t}\in\mathbb{R}^{D\times D}$ is the state matrix which is recursively updated along time. The parallel representation Eq. (1) is suitable for efficient training, while the recurrent representation Eq. (2) has a linear temporal complexity and can be used for inference. In between, [30] also provides a chuckwise recurrent representation, which facilitates the training with long sequences. Within each chunk, computation follows the parallel representation as shown in Eq. (1). While state recurrence is performed across chunks using the recurrent representation in Eq. (2). Specifically, let $B$ and $i$ denote the length and index of chunk, respectively. The input sequence, key, query and value of the $i$-th (non-overlapped) chunk are denoted as $X_{[i]},K_{[i]},Q_{[i]},V_{[i]}\in\mathbb{R}^{B\times D}$, respectively. The chunk-wise state matrix $R_{i}\in\mathbb{R}^{D\times D}$ can be recursively computed as:
$R_{i}=K_{[i]}^{\top}(V_{[i]}\odot\zeta)+\gamma^{B}R_{i-1},\ \zeta_{mn}=\gamma^{B-m-1},$ (3)
and then the chunk-wise Retention can be computed as
$\text{Retention}(X_{[i]})=$ $\underbrace{(Q_{[i]}K_{[i]}^{\top}\odot\Gamma)V_{[i]}}_{\text{Inner-chunk}}+\underbrace{(Q_{[i]}R_{i-1})\odot\xi}_{\text{Cross-chunk}},$
$\xi_{mn}=\gamma^{m+1},$ (4)
where $\Gamma\in\mathbb{R}^{B\times B}$ and $\gamma$ are similarly defined as in Eq. (1). This design allows Retention to maintain linear memory complexity for processing long sequences.
Removing softmax could degrade the numerical scale stability. To compensate for this, three extra normalization factors are applied to the attention scores of $QK^{\top}\odot\Gamma$ in Eq. (1) and a group normalization layer is added right after the retention layer, please refer to [30] for more details.
Retention also uses multiple heads, and multi-scale retention assigns different $\gamma$ for each head as
@@
where $h$ denotes the number of Retention heads. $\gamma$ controls the forgetting rate of past information. In addition, after the multi-head retention and the group normalization layer, an extra swish gate [45] is applied to increase the non-linearity of retention layers.
Speaker diarization learns voiceprint of speakers, which are time-invariant in a very long time period, e.g. in the entire audio stream to be processed. In the proposed model, to keep the consistency of speaker embedding along the entire audio stream, we set $\gamma=1$ to not have information decay.
@@ III-B3 Online Attractor Decoder
Decoder Input. At each frame, the speaker embedding is fed into the decoder to generate/update the corresponding attractors. All the $S+2$ attractors share the same repeated embedding as their input source, and the decoder determines whether the embedding is used for attractor updating. To distinguish the input to multiple attractors from each other, we first repeat the speaker embedding $S+2$ times, which are then concatenated (in the feature dimension) with a positional embedding (PE) [47] along the speaker/attractor index (speaker-index PE):
$\begin{split}\text{Speaker-Index PE}(s,2d)&=sin(s/10000^{2d/D)},\\ \text{Speaker-Index PE}(s,2d+1)&=cos(s/10000^{2d/D}),\end{split}$ (6)
where $s$ indexes the speakers/attractors, and $d$ indexes the feature dimension. Note that the same speaker-index PE is utilized for all time frames. A linear layer is then applied to restore the feature dimension $D$. Thus, the input dimension of the attractor decoder is $T\times(S+2)\times D$}.
Decoder Architecture. To derive the attractor $\boldsymbol{a}_{s,t}$ of the $s$-th speaker at frame $t$, besides the speaker embedding extracted at frame $t$, two more information sources need to be considered:
  • The attractor of previous frames for the same speaker, i.e., $\boldsymbol{a}_{s,t^{\prime}},t^{\prime}<t$, which helps maintain the attractor’s consistency of a specific speaker over time. To refer to the extracted attractor of previous frames, recursive computation is normally required, which however causes implementation difficulties of parallel computation during training. Instead, we design a non-recursive scheme by applying a temporal Retention module in the hidden space of attractors, which is also proved to be effective for retrieving the attractor information of previous frames. Specifically, let $H\in\mathbb{R}^{T\times(S+2)\times D}$ denote the hidden units of attractors. The temporal Retention module processes the hidden sequence of each speaker/attractor independently, and the attractor axis is taken into the batch dimension:
    $\text{Temporal Retention}(H[:,s,:]).$ (7)
    For the same reason as stated above, in temporal Retention, $\gamma$ is set to 1 and the positional embedding terms are removed.
  • The attractor of other speakers at frame $t$, i.e., $\boldsymbol{a}_{s^{\prime},t},s^{\prime}\neq s$. By leveraging information from other attractors, the distance between attractors can be enlarged to better distinguish different speakers. To this end, we develop a cross-attractor self-attention module which processes each time frame independently, following the standard multi-head self-attention mechanism [47]:
    $\text{Cross-Attractor Self-attention}(H[t,:,:]).$ (8)
After, a feed-forward module is employed. Each of the three modules incorporates a residual connection and layer normalization. $M$ decoder blocks are cascaded.
Decoder Output. The decoder output vectors are then L2-norm normalized to have unit length, serving as the final attractors $\boldsymbol{A}_{t}=(\boldsymbol{a}_{0,t},\boldsymbol{a}_{1,t},\ldots,\boldsymbol{a}_{S,t},\boldsymbol{a}_{\text{ter},t})$. Speaker activities are calculated by taking the inner product between the frame-wise attractors and the speaker embedding $\boldsymbol{e}_{t}$ as:
$\hat{\boldsymbol{y}}_{t}=\sigma(\boldsymbol{A}_{t}^{\top}\boldsymbol{e}_{t})\in(0,1)^{S+2},$ (9)
@@ III-C Training Loss
where $\mathbb{S}=\{0,1,\ldots,S,{\text{ter}}\}$. The diarization loss trains the network to: i) maintain the consistency of speaker embedding (and also its corresponding attractor) extracted at different frames for the same speaker, and disentangle out speech content information; ii) better discriminate speaker embeddings (and also the corresponding attractors) for different speakers, by leveraging such as their different vocal tract attributes (e.g., length and shape), vocal fold properties (e.g., fundamental frequency F0), and speaking styles (e.g., prosody including rhythm and intensity)[48, 49]; and iii) to be able to handle overlapped speakers in terms of both speaker embedding learning and attractor learning, by performing kind of multi-class classification.
Note that both the embeddings and attractors are estimated within each recording and are only used for speaker diarization of that recording. Embeddings (and also attractors) could be different for the same speaker when appearing in different recordings. A similar observation is discussed in [13].
Embedding similarity loss. The clustering effect of the frame-wise speaker embeddings has a salient impact on diarization performance. In addition to the indirect training of speaker embeddings made by the diarization loss. To regulate the distribution of these embeddings, we introduce an extra direct embedding similarity loss. The cosine similarity between each pair of embeddings ($\langle\boldsymbol{e}_{j},\boldsymbol{e}_{k}\rangle,j,k\in\{1,\ldots,T\}$) is constrained to be consistent with the cosine similarity between their corresponding diarization labels ($\langle\boldsymbol{y}_{j},\boldsymbol{y}_{k}\rangle,j,k\in\{1,\ldots,T\}$). The mean square error between the two cosine similarities is minimized as
$\mathcal{L}_{e}=\frac{1}{T(T-1)/2}\sum_{j=1}^{T}\sum_{k=j+1}^{T}(\langle\boldsymbol{e}_{j},\boldsymbol{e}_{k}\rangle-\langle\boldsymbol{y}_{j},\boldsymbol{y}_{k}\rangle)^{2},$ (11)
where $\langle\cdot,\cdot\rangle$ denotes cosine similarity. It should be noted that the embedding similarity loss is suitable for various situations, including two frames from the same, different, and overlapped speakers. The higher the overlap between two frames, the smaller the angle between their corresponding embedding vectors should be.
@@ III-E Inference
At inference, the proposed model processes audio streams frame by frame, with the recurrent paradigm of Retention and a linear computational complexity w.r.t frames. The maximum number of speakers $S$ is set to the maximum number appeared in the training set. Regardless of the actual number of speakers, the network always outputs activity probabilities for $S+2$ attractors. Since the first attractor represents non-speech, speech activity are evaluated starting from the second attractor. When the actual number of speakers is smaller than the preset $S$, it can be automatically detected by the speaker termination marker.
@@ IV-A1 Simulated Mixtures
[…]